Two resistors R_1 and R_2 are connected in series
Answer #1

Net resistance of circuit will be

R=R_1+R_2=10+5=15

Hence current will be

I=\frac{V}{R}=\frac{15}{15}=1

Hence power dissipated at R_2 will be

H=i^2R_2=1^2\times5=5$ J$